OpenStudy (anonymous):

can anyone explain to me how 25(4a^2-4ab+b^2) = 25(2a-b)^2? thankyouu!

OpenStudy (anonymous):

\[(2a-b)^2=(2a)^2-2(2a)(b)-b^2 = 4a^2-4ab-b^2\]

OpenStudy (anonymous):

25(4a^2-4ab+b^2) = 25[ (2a)^2-2(2a)(b)+b^2]=25[2a-b]^2

OpenStudy (anonymous):

thank-you